a proton and a duteron are both accelerated through the same potential difference and enter in a magnetic field perpendicular to the direction of the field if the deutron follows a path of radius r assuming the neutron and proton masses are nearly equal the radius of a proton path will be?

Grade:12

1 Answers

Arun
5 years ago

For proton (r)p= X cm (let)

For deutron (r)d = [2(m)d(k)d]^1/2/eB = [4(m)p(k)p]^1/2/eB = 2^1/2X or 1.41x cm

 

So the ratio is 1:2^1/2